iuytyuioiuytyuiop:

If (x + 5) - 2(4x - 1) = 0, what is the value of x?

iuytyuioiuytyuiop:

can someone teach me step by step

supie:

So basically we are solving for x do you know the first step?

iuytyuioiuytyuiop:

no

imHoodKingTae:

What u need help wit??

callmeizzy1234:

okay first thing to do is go left to right

iuytyuioiuytyuiop:

the first step?

darkknight:

(x + 5) - 2(4x - 1) = 0, first distribute -2(4x-1) into -8x+2

callmeizzy1234:

5+x is your first part of the problem now th odesnt eaqual 5x cause there not like terms so you levae it alone

imHoodKingTae:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @callmeizzy1234 5+x is your first part of the problem now th odesnt eaqual 5x cause there not like terms so you levae it alone \(\color{#0cbb34}{\text{End of Quote}}\) Mhm

callmeizzy1234:

now distribue you 4

darkknight:

(x + 5) - 2(4x - 1) = 0, combine x+5 with -8x+2 and get -7x + 7 =0 solve for x from here

callmeizzy1234:

-2 not 4

callmeizzy1234:

so -2 x 4 and -2 x -1

iuytyuioiuytyuiop:

you guys are giving me different ways to solve but i only need one

darkknight:

@iuytyuioiuytyuiop listen to me ig

iuytyuioiuytyuiop:

Ok

callmeizzy1234:

now and negitve plus a positve is a negitive and a negitive plus a negitve is a positive

iuytyuioiuytyuiop:

where did you get -8x+2

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight (x + 5) - 2(4x - 1) = 0, first distribute -2(4x-1) into -8x+2 \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight (x + 5) - 2(4x - 1) = 0, combine x+5 with -8x+2 and get -7x + 7 =0 solve for x from here \(\color{#0cbb34}{\text{End of Quote}}\)

iuytyuioiuytyuiop:

ok guys one last thing

iuytyuioiuytyuiop:

what topic is this called?

callmeizzy1234:

you mutilpy -2 x 4 and -2 x -1

jahniya59:

so if te 2 number is parthesses you know to multiply going on between them 1 is equalled to 0

iuytyuioiuytyuiop:

so i can look it up and practice more similar problems

jahniya59:

then

iuytyuioiuytyuiop:

@supie or @darkknight do you know what the topic is called?

callmeizzy1234:

now combine like trems so with what i have already said you should have 5+x-8x+2

callmeizzy1234:

so now you combine 8x and x to get 9 x

iuytyuioiuytyuiop:

my life is on the LINE

iuytyuioiuytyuiop:

HELP

iuytyuioiuytyuiop:

what topic is this called?

callmeizzy1234:

and you get 5+9x+2

iuytyuioiuytyuiop:

is it a polynomial or something else?

callmeizzy1234:

now you need to get the x alone so you - 9 to both sides so you get -4+9x

jahniya59:

there both 0 bec any number times it self is 0 so x-5=0 and x+2=0 so on the x-5 it o you have to + on both sides so it = 5

callmeizzy1234:

then divide by nine

callmeizzy1234:

-4/9 is what and thats your answer

iuytyuioiuytyuiop:

thank you

callmeizzy1234:

np

iuytyuioiuytyuiop:

wait

callmeizzy1234:

yea

jahniya59:

in on the x-2=0 you have to - on both sides that is negative is =5,-2

iuytyuioiuytyuiop:

can you tell me what this topic is so i canpractice similar problems?

callmeizzy1234:

mathmatics

jahniya59:

it mostly just multply if you then you have to add on both sides

iuytyuioiuytyuiop:

i mean like is it called polynomials or

iuytyuioiuytyuiop:

so i can look it up on khan academy

callmeizzy1234:

i think its polynomials

jahniya59:

yes

iuytyuioiuytyuiop:

ok

callmeizzy1234:

you can also do ixl

iuytyuioiuytyuiop:

ok

jahniya59:

it expressions

callmeizzy1234:

it is very helpful tool as well

iuytyuioiuytyuiop:

its expressions?!

jahniya59:

with absolute value

jahniya59:

yes

iuytyuioiuytyuiop:

is he right?

jahniya59:

yes

iuytyuioiuytyuiop:

suple?

jahniya59:

but it real expression in the right word

iuytyuioiuytyuiop:

ohhhhhhhhh ok ill look up real exppression on khan academy

iuytyuioiuytyuiop:

can anyone else agree?

iuytyuioiuytyuiop:

@supie

jahniya59:

An expression is a sentence with a minimum of two numbers and at least one math operation. This math operation can be addition, subtraction, multiplication, and division. The structure of an expression is: Expression = (Number, Math Operator, Number)

iuytyuioiuytyuiop:

ok

jahniya59:

np